OpenStudy (anonymous):

how do you simplify -3+-√63/-4 farther? I'm stuck.

mathslover (mathslover):

Did u mean this : \[\huge{\frac{-3\pm\sqrt{63}}{-4}}\]

OpenStudy (anonymous):

yes.

OpenStudy (turingtest):

I think the technical term for this kind of expression is "ugly" no I don't think you can simplify anything

OpenStudy (anonymous):

the answer is left as it is....its irrational and u cannot and need not simplify it.

mathslover (mathslover):

U can calculate this : But u can not simplify this

OpenStudy (anonymous):

idk wtf this guy teached me then?

OpenStudy (anonymous):

the answers are x=3+-i√29/2 < and then the same but just -3 and then x=3+-i√23 and then the same just with -3 thses are my choies

OpenStudy (anonymous):

\[-3/-4\pm \sqrt{63}/-4; 3/4\pm \sqrt{63/16}; 0.75 \pm \sqrt{3.9375}\]

OpenStudy (helder_edwin):

the final result would be \[\frac{3\pm 3\sqrt{7}}{4}\]

mathslover (mathslover):

ok clear i think we can simplify this further : \[\huge{\frac{-3\pm\sqrt{7*9}}{-4}}\] \[\huge{\frac{-3\pm3\sqrt{7}}{-4}}\]

OpenStudy (anonymous):

I know and that's what I dnt get cuz that's not one of my choices

OpenStudy (anonymous):

oh but u can write the square root as the one helder_edwin mentioned 63=9*7 so square root of 9 simplifies to 3

mathslover (mathslover):

Can u give the whole question to us please ?

OpenStudy (anonymous):

Solve x^2 – 3x = –8

OpenStudy (helder_edwin):

the equation would be then x^2-3x+8=0.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yea then u made a mistake the answer is ur first choice....remember solution to quadratic eqn contains the term b^2-4ac??

OpenStudy (helder_edwin):

this solves into \[ \frac{3\pm \sqrt{9-4\cdot 1\cdot 8}}{2}=\frac{3\pm\sqrt{-23}}{2}=\frac{3\pm \sqrt{23}i}{2} \]

OpenStudy (anonymous):

wait which one is it, you guys gave 2 different answers?

OpenStudy (anonymous):

the above is right

OpenStudy (anonymous):

okay thank you guys very much, you were very helpful

OpenStudy (helder_edwin):

you are welcome

OpenStudy (anonymous):

:D